OpenStudy (anonymous):

Find a general formula for f^(n)(x) if f(x) = 8xe^−x

zepdrix (zepdrix):

Hey cassie :) Have you tried taking a few derivatives? You might see a pattern!

zepdrix (zepdrix):

\[\Large\rm f(x)=8xe^{-x}\]\[\Large\rm f'(x)=8e^{-x}-8x e^{-x}\]The trick is to make sure you're factoring each time.\[\Large\rm f'(x)=8e^{-x}(1-x)\]

geerky42 (geerky42):

Here's how I see the pattern: \[f(x)=8xe^{-x} \\~\\f'(x) = 8e^{-x}-8xe^{-x} = 8e^{-x}-f(x)\\~\\f''(x) = -8e^{-x}-f'(x) = -8e^{-x} - [8e^{-x}-f(x)] = 2(-8e^{-x})+f(x)\\~\\f'''(x)=8e^{-x}-f''(x) = 8e^{-x}-[2(-8e^{-x})+f(x)] = 3(8e^{-x})-f(x)\] To make pattern more clear: \[f(x)=0(-8e^{-x}) + f(x)\\~\\f'(x) = 1(8e^{-x})-f(x)\\~\\f''(x) =2(-8e^{-x})+f(x)\\~\\f'''(x)= 3(8e^{-x})-f(x)\\~\\f''''(x) = 4(-8e^{-x})+f(x)\]See the pattern, right?

OpenStudy (anonymous):

ah yes thank you I was having difficulty finding the pattern!

geerky42 (geerky42):

Welcome :)